OpenStudy (anonymous):

The length of one leg of a right triangle is 8ft. The length of the hypotenuse is 4ft longer than the other leg. Find the length of the hypotenuse and the other leg.

OpenStudy (anonymous):

The length of the leg other leg is 15 ft. and the length of the hypotenuse is 17 ft.. You represent the legs as the following variables: other leg = x therefore hypotenuse is x+2 since it is 2ft longer. Plug this into pythagorean theorem and solve for x : (x+2)^2 = x^2 + 8^2. X^2+4x+4 = X^2+64 4x=60 so 60/4 =15 ( other leg) so 15 +2= hypotenuse. . * the answer in the post above (colinsnewid), it cant be 8+10 because its two feet bigger than the other unknown leg, not the leg that is already given.